OpenStudy (anonymous):

f'(x)=e^-x^2 f(5)=1 Find f(2). Thanks!

OpenStudy (anonymous):

does this say \[f(x)=e^{-\frac{x^2}{2}}\]??

OpenStudy (solomonzelman):

integrate, and based on f(5)=1 find C. then find f(2)

OpenStudy (solomonzelman):

are you sure that is the derivative ?

OpenStudy (anonymous):

positive

OpenStudy (solomonzelman):

then f(5)\(\large\color{slate}{\displaystyle\ne}\)1

OpenStudy (anonymous):

how do you integrate with the e? and the end of your post is coming out in code... not sure what its supposed to be

OpenStudy (solomonzelman):

if that is the derivative, \(\large\color{slate}{f'(x)=e^{-\frac{x^2}{2}}}\). then you can use the error function, or maybe a power series.......

OpenStudy (solomonzelman):

you will not find a simple result. (I don't think so)

OpenStudy (solomonzelman):

If it was something like \(\large\color{slate}{f'(x)=xe^{-\frac{x^2}{2}}}\) that would be much neasier.

OpenStudy (solomonzelman):

*easier.

OpenStudy (anonymous):

What is the (large/color(slate) thing?

OpenStudy (solomonzelman):

oh, you don't see any cool math symbols, just some random codes?

OpenStudy (anonymous):

yup

OpenStudy (solomonzelman):

that is then a latex-processing error.

OpenStudy (solomonzelman):

I will draw what I said.